OpenStudy (anonymous):

integrate 36(4x^2-9)^-1

OpenStudy (zugzwang):

\[\int\limits_{}^{}\frac{36}{4x^{2}-9}dx\] is it?

OpenStudy (zugzwang):

Just use the fact that \[\frac{36}{4x^{2}-9}=\frac{6}{2x-3}-\frac{6}{2x+3}\]

OpenStudy (anonymous):

yes , then how to integrate it ? one by one or what?

OpenStudy (zugzwang):

Well, the integral sign distributes over addition (and over subtraction) And use the fact that for any nonzero a, b, and c \[\int\limits\limits_{}^{}\frac{c}{ax+b}dx=\frac{c}{a}\ln|ax+b|+K, \ \ \ K \in \mathbb{R}\]

OpenStudy (anonymous):

what formula is that ? never seen it .

OpenStudy (zugzwang):

\[\int\limits_{}^{}\frac{c}{ax+b}dx\]Let u=ax+b then du = adx \[\int\limits\limits_{}^{}\frac{c}{ax+b}dx=c\int\limits\limits_{}^{}\frac{1}{ax+b}dx=\frac{c}{a}\int\limits\limits_{}^{}\frac{1}{u}du\] \[=\frac{c}{a}\ln \left| u \right|+K\] \[=\frac{c}{a}\ln \left| ax+b \right|+K\]

OpenStudy (anonymous):

okay ! thanks :)